# The Kernel Manifold: A Geometric Approach to Gaussian Process Model Selection

### v1

@@ 3.4 Transformations for Euclidean Embeddability
Figure 7: MDS reconstruction error as a function of embedding dimension for the log-transformed Hellinger distance matrix. Although log-warping improves curvature and stabilizes the Gram spectrum, the induced geometry remains high-dimensional: the reconstruction error decays slowly and only approaches zero at large embedding dimensions ($k\approx 80$$100$), indicating that the transformed distance matrix has high intrinsic rank.
@@ 4.2 Comparison with LLM-guided Genetic Algorithm (LLM-GA) search method
4.3 Case Study: Melt-Pool Geometry Data with Process variability in AM process
@@
4.4 Bayesian Optimization with the Optimized Kernel
@@ 5 Conclusion
The code used to generate the results in this study is openly available at https://github.com/shafiqmme/Kernels-BO. A permanent, versioned archive of the code is available via Zenodo at https://doi.org/10.5281/zenodo.18189228.

### v2

@@ 3.4 Transformations for Euclidean Embeddability
Figure 7: MDS reconstruction error as a function of embedding dimension for the log-transformed Hellinger distance matrix. Although log-warping improves curvature and stabilizes the Gram spectrum, the induced geometry remains high-dimensional: the reconstruction error decays slowly and only approaches zero at large embedding dimensions ($k\approx 80$100100), indicating that the transformed distance matrix has high intrinsic rank.
@@ 4.2 Comparison with LLM-guided Genetic Algorithm (LLM-GA) search method
4.3 Budgetary impact of inline kernel optimization during BO
While figure 17 reports the absolute computational overhead of kernel optimization, its practical impact must be interpreted relative to the cost of objective evaluations in real Bayesian optimization workflows. In many scientific and engineering applications, a single BO iteration corresponds to an expensive experiment or simulation, often ranging from several minutes to hours or days.
For example, we can consider a realistic setting in which BO is initialized with 50 observations from a black-box objective. Performing kernel optimization on this static dataset incurs an upfront cost of approximately 1–2 minutes. Subsequent BO iterations may each require on the order of 10 minutes (or more) per evaluation. After 10 BO iterations, the dataset grows to 60 points, and re-optimizing the kernel again incurs a comparable marginal cost. In this regime, kernel optimization contributes roughly 2 minutes of overhead relative to approximately 100 minutes spent on objective evaluations, corresponding to a $\sim$2% increase in total runtime.
As the cost of individual BO evaluations increases, e.g., to 30 minutes, several hours, or even days of experiments, the relative overhead introduced by kernel optimization decreases rapidly and becomes effectively negligible, as illustrated in Fig. 18. Consequently, inline kernel optimization does not meaningfully inflate the total BO budget and is well justified in settings where surrogate fidelity is critical to sample efficiency.
Figure 18: Illustrative scaling of the relative computational overhead introduced by inline kernel optimization as a function of the cost per Bayesian optimization (BO) iteration. The curve shows the fraction of total runtime attributable to kernel optimization under a simple budget model with periodic kernel re-optimization as the dataset grows. As the cost of individual BO evaluations increases—from minutes to hours or days—the relative overhead from kernel optimization decays rapidly and becomes negligible. This figure is schematic and intended to illustrate scaling behavior rather than report additional benchmark measurements.
4.4 Case Study 1: Melt-Pool Geometry Data with Process variability in AM process
@@
4.5 Case Study 2: Bayesian Optimization with the Optimized Kernel
@@ 5 Conclusion
The code used to generate the results in this study is openly available at https://github.com/shafiqmme/Kernels-BO. A permanent, versioned archive of the code is available via Zenodo at https://doi.org/10.5281/zenodo.18234227.